Two plane waves of light with wavelength 500 nm are incident on a screen. If the

spatial resolution of the detector is 0.1 mm, what is the maximum angle between

the two waves for which we can resolve the fringes. Assume we can resolve fringes if

the distance between the maxima and minima is equal to the detector resolution.


1
Expert's answer
2021-11-21T17:27:06-0500

"\sin \theta=\theta=\frac{2\lambda}d=0.57°."